OpenStudy (anonymous):

work done by performing a reversible, isothermal compression of 2 moles of an ideal gas at T = 300K from V1 = 2L to V2 = 1L. i get -3456 j/mol, but im not overly confident Thanks

OpenStudy (anonymous):

@styx_83 How did you arrive at your answer?

OpenStudy (anonymous):

\[w=-\int\limits_{V1}^{V2}PdV \] \[pV=nRT\] \[w=-\int\limits_{V1}^{V2}\frac{ nRT }{ V } dV =-nRT ln \frac {V_2}{V_1} \] \[w =-2\cdot 8.314 \cdot 300 \cdot \ln 2 = - 3.458 kJ/mol\]

OpenStudy (anonymous):

You have calculated the total work by multiplying by 2. \[w=W/n= 3458/2 = 1729j/mole\] I assume you are using gram molecular wgt also

OpenStudy (anonymous):

@gleem Yes, I have made a mistake. What I have written is actually for 2 mols.